Comments: Following the switch from Submode 6 (which was delayed from yesterday because of an antenna problem) the active region campaign ran successfully from 17:00 - 21:30. CDS had switched their observing time to earlier in the day, while Sac Peak observed at the same time as EIT.

The 195 Å campaign continues, with half-resolution images every 17 minutes instead of full-resolution images at approximately a 12-minute cadence.

Calibration exercises will continue through the week in anticipation of the bakeout which is currently scheduled to begin on 18-July-1997. BJT